# WatME: Towards Lossless Watermarking Through Lexical Redundancy

### v2

@@ Abstract
Text watermarking has emerged as an important technique for detecting machine-generated text. However, existing methods generally use arbitrary vocabulary partitioning during decoding, which results in the absence of appropriate words during the response generation and disrupts the language model’s expressiveness, thus severely degrading the quality of text response. To address these issues, we introduce a novel approach, Watermarking with Mutual Exclusion (WatME). Specifically, by leveraging linguistic prior knowledge of inherent lexical redundancy, WatME can dynamically optimize the usage of available vocabulary during the decoding process of language models. It employs a mutually exclusive rule to manage this redundancy, avoiding situations where appropriate words are unavailable and maintaining the expressive power of large language models (LLMs). We present theoretical analysis and empirical evidence demonstrating that WatME substantially preserves the text generation ability of LLMs while maintaining watermark detectability. Specifically, we investigate watermarking’s impact on the emergent abilities of LLMs, including knowledge recall and logical reasoning. Our comprehensive experiments confirm that WatME consistently outperforms existing methods in retaining these crucial capabilities of LLMs. Our code will be released to facilitate future research.
@@ 1 Introduction
The advent of large language models  Ouyang et al. (2022); OpenAI (2023a) with human-level generative capabilities presents tremendous opportunities across diverse domains. However, their ability to synthesize high-quality text also raises widespread concerns about potential misuse, including the dissemination of misinformation  Zellers et al. (2019) and the facilitation of academic dishonesty  Stokel-Walker (2022). This necessitates developing techniques to reliably attribute generated text to AI systems.
@@
Existing approaches typically fall into two main paradigms. The first type attempts to distinguish machine-generated text by hunting for inductive statistical or linguistic patterns (Gehrmann et al., 2019; Mitchell et al., 2023; Zellers et al., 2019; OpenAI, 2023b), employing methods that span from basic manual feature engineering to the intricate training of complex classifiers. However, as generative models continue improving, their outputs increasingly resemble the pattern of human writing, rendering statistical detectors ineffective (Dou et al., 2022; Sadasivan et al., 2023; Chakraborty et al., 2023). The second paradigm promotes a more proactive approach, advocating for direct intervention in the generative process to actively watermark model outputs (Kirchenbauer et al., 2023; Christ et al., 2023; Zhao et al., 2023). This strategy embeds identifiable fingerprints within machine-generated text, enabling provenance verification. As LLMs’ capabilities continue to grow, this approach is more effective in detecting LLM-generated text Sadasivan et al. (2023). However, introducing watermarks during text generation can significantly impact output quality, which has been a consistent challenge for model developers - how to effectively watermark while preserving text quality.
@@
In this paper, we introduce a novel approach to text watermarking by leveraging engineered lexical redundancy during the decoding phase of language generation. Our method utilizes the comprehensive set of tokens available to a language model, clustering them based on overlapping semantic or syntactic functionalities to create sets of interchangeable tokens. This process simulates redundancy within the lexical space, akin to the surplus pixels in images that facilitate watermarking in multimodal data Nikolaidis and Pitas (1999); Samuel and Penzhorn (2004). The motivation for this strategy arises from the challenge of applying traditional watermarking techniques to textual data. In contrast to the inherent redundancy found in images, the discrete and succinct nature of textual data offers little to no native redundancy, making it challenging to exploit redundancy in the textual space zho1u-etal-2021-defense; He et al. (2022). By engineering lexical redundancy, our method not only surmounts the limitations imposed by the inherent properties of natural language but also paves the way for secure and efficient text watermarking.
@@ 2 Related Work
Early works on AI-generated text detection develop post-hoc detection methods to analyze machine-generated text by treating the problem as a binary classification task (OpenAI, 2019; Jawahar et al., 2020; Mitchell et al., 2023). For instance, OpenAI has fine-tuned RoBERTa (Liu et al., 2019) to distinguish between human and GPT-2 generated texts (OpenAI, 2019). However, existing detectors are found to be fragile against adversarial attacks (Wolff, 2020) and biased towards non-native English writers (Liang et al., 2023). Moreover, as LLMs continue to advance, their generated outputs more closely resemble human-written text, rendering these methods progressively less effective.
On the other side, watermarking, traditionally a copyright marking method Adi et al. (2018); Rouhani et al. (2018), involves developers, users, and regulatory entities. Developers choose an algorithm to subtly embed hidden modifications into data, which can be altered during user transmission. Regulatory bodies can later extract this information to trace and regulate AI-generated content (Atallah et al., 2001; Wilson et al., 2014; Hacker et al., 2023). In the context of natural languages, watermarking typically involves modifying content or structure. For example, rule-based methods (Stefan et al., 2000) or carefully designed neural encoders (Yang et al., 2022; Ueoka et al., 2021) encrypt messages into text, which are then extracted using the corresponding rules and neural decoder. The discrete nature of natural language, however, presents a considerable challenge to this approach, as modifications can unintentionally degrade text quality or alter its intended meaning.
For the detection of LLM-generated texts, a pioneering watermarking technique (Kirchenbauer et al., 2023) partitions tokens into ’green’ and ’red’ lists, biases output distribution towards ’green’ tokens, and creates patterns that are detectable yet imperceptible to humans. Nevertheless, while yielding promising detection results, these methods may still degrade textual quality and be vulnerable to the paraphrase attack. Current efforts Christ et al. (2023); Fernandez et al. (2023); Zhao et al. (2023) in this field aim to develop more robust watermarking methods capable of defending various user attacks.
@@ 3 Method
In this section, we begin by providing a summary of the preliminaries related to text watermarking. Subsequently, we delve into an investigation of redundancy in the lexical space and demonstrate how this redundancy can be leveraged to develop a watermark algorithm that achieves a higher degree of losslessness for large language models. Finally, we employ mathematical analysis to elucidate the benefits of our proposed method.
@@ 3.3 WatME: Exploit the Lexical Redundancy
We then detail the WatME encoding process, outlined in Alg. 1, which employs a two-step partitioning process to form a green and red list. The first partition occurs within the redundant lexical clusters $C$ that we have identified within $S$, while the second takes place among the remaining part in the vocabulary denoted as $\mathcal{V}\setminus\mathcal{S}$. We use $\gamma$ to determine the number of tokens from the mined clusters that are allocated to the green list $G_{t}^{\prime}$ in the first partition. The remaining tokens, based on the principle of mutual exclusivity, are assigned to the red team $R_{t}^{\prime}$. The second partition continues to allocate words to the green list $G_{t}$ from the remaining vocabulary until the combined size of the green teams from both steps reaches the predefined limit, $\gamma$. The rest of the process follows the steps outlined in the vanilla watemarking of Alg. 2.
@@ 5.1 Experimental Setups
[…]
@@
[…]
@@
We compared our model with four established baselines. First, KGW-Mark Kirchenbauer et al. (2023), which categorizes teams into ’red’ and ’green’ to facilitate detection. Second, Gumbel-Mark Kuditipudi et al. (2023), which employs a Gumbel-Softmax distribution to introduce stochasticity into the watermarking process. Third, Unbiased-Mark Hu et al. (2024), which implements reweighting techniques to maintain the expected output distribution of the LLM during watermarking. Lastly, Provable-Mark Zhao et al. (2023), which uses a fixed hash key during watermarking to achieve provably better performance.
@@ 5.2 Main Results
Despite the trade-off between text quality and detection performance, WatME’s detection efficacy matched that of the Vanilla Watermark while also enhancing model capabilities, as evidenced by similar AUROC scores—suggesting our algorithm attained a better equilibrium than the baseline. In contrast, the Gumbel-Mark method noticeably compromised detection performance, particularly in aligned models and when generating short responses (TruthfulQA). Additionally, more performance results under different watermark strength are presented in Discussion 6.3.
@@ 6.1 Analysis of Clustering Methods
To analyse redundant clusters from diverse methods, we set evaluation criteria to ensure analytical rigour. These criteria spanned semantic consistency, contextual appropriateness, and grammatical consistency, which are essential aspects for cluster quality. Two annotators used a rating scale of 0, 1, 2 to annotate the clusters. A score of ’2’ indicated high or ideal consistency, ’1’ denoted moderate or usable consistency, and ’0’ identified low or unusable consistency within a cluster. The kappa value for the annotations is 0.657. Figure 2(a) shows both methods met usability, but fell short of ideal effectiveness. The dictionary approach was superior in semantic coherence due to its utilization of lexical databases. Conversely, the prompting method outperformed in contextual and grammatical consistency, reflecting the varied linguistic corpus training of LLMs. This suggests the potential benefits of a combined approach, a topic reserved for future research.
@@ 6.2 Robustness Against Attacks
Within the scope of watermark robustness against common rewriting attacks, our study evaluated the resilience of the proposed WatME method compared to baseline watermarking techniques. In a simulated black-box attack scenario, where attackers were blind to the watermark encryption algorithm, we assessed the integrity of watermarks after random substitutions of text tokens. Utilizing a sample of 200 examples from the GSM8k dataset, the analysis, illustrated in Figure 2(b), demonstrated that WatME consistently outperformed vanilla method in detection robustness across a spectrum of replacement ratios.
@@ 6.3 Performance Trade-offs at different Delta
The efficacy of Watermark is influenced by the hyperparameter, Delta, which controls the watermark strength. An increase in Delta facilitates easier watermark detection but at the cost of severer impact on the LLMs. We analyse on the TruthfulQA and GSM8K datasets. Figure 3 shows WatME consistently achieved a more favorable balance between watermark robustness and LLM performance across various Delta settings, surpassing Vanilla Watermark. Notably, the performance curves of WatME are strictly better than that of Vanilla, indicating that at equivalent watermark strengths, WatME always maintains superior performance compared to Vanilla Watermark.
@@ 6.4 Aligned vs Unaligned Models
Our examination of the response sensitivity to watermarking in aligned and unaligned models involved analyzing their output distributions on the TruthfulQA and GSM8K datasets. We computed the average entropy for token in the generated answers and found that aligned models exhibit markedly lower entropy, suggesting more deterministic response patterns, as illustrated in Figure 4. This pronounced certainty in aligned models’ outputs presents a challenge for watermarking because of the limited variability that is essential for effective watermark encoding.
@@ 7 Conclusion
Lastly, the complexity of selecting contextually appropriate synonyms for text watermarking necessitates advanced synonym-handling processes. Our approach incorporates syntactic elements and leverages LLMs to ensure the contextual appropriateness of synonyms. However, the inherent complexity of this task, which requires predefined rules for on-the-fly watermarking, represents a limitation that warrants further discussion and exploration in future work.
@@ Appendix F Examples of Redundant Clusters
We present some examples of mined clusters at 6.
Figure 6: Examples of Redundant Clusters.
Method GSM8K TruthfulQA
He et al. (2022) 0.5463 0.5825
WatME 0.9128 0.8659
Table 2: Performance comparison of different methods on GSM8K and TruthfulQA benchmarks.
Appendix G Comparison with Model Theft Prevention Watermarking
Our approach, WatME, differs significantly from the method presented in He et al. (2022), which is aimed at model theft prevention. While He et al. (2022) focuses on tracing model theft, WatME concentrates on text watermarking to identify AI-generated text. Additionally, WatME is designed to be effective with a single sample of text, contrasting the necessity for large volumes of text required by He et al. (2022) for effective model theft detection. Furthermore, our method does not rely on the presence of a trigger word, as He et al. (2022)’s method does, which may not be present in short texts. This makes WatME more versatile and applicable to a broader range of text lengths and types, embedding watermarks without the need for specific trigger words or phrases.
As shown in 2, We experimented with the approach suggested in He et al. (2022) for text watermarking but encountered failures in detection performance (AUROC).

### v3

@@ Abstract
Text watermarking has emerged as a pivotal technique for identifying machine-generated text. However, existing methods often rely on arbitrary vocabulary partitioning during decoding to embed watermarks, which compromises the availability of suitable tokens and significantly degrades the quality of responses. This study assesses the impact of watermarking on different capabilities of large language models (LLMs) from a cognitive science lens. Our finding highlights a significant disparity; knowledge recall and logical reasoning are more adversely affected than language generation. These results suggest a more profound effect of watermarking on LLMs than previously understood. To address these challenges, we introduce Watermarking with Mutual Exclusion (WatME), a novel approach leveraging linguistic prior knowledge of inherent lexical redundancy in LLM vocabularies to seamlessly integrate watermarks. Specifically, WatME dynamically optimizes token usage during the decoding process by applying a mutually exclusive rule to the identified lexical redundancies. This strategy effectively prevents the unavailability of appropriate tokens and preserves the expressive power of LLMs. We provide both theoretical analysis and empirical evidence showing that WatME effectively preserves the diverse capabilities of LLMs while ensuring watermark detectability. Our code will be released to facilitate future research. via https://github.com/ChanLiang/WatME.
@@ 1 Introduction
The advent of large language models Ouyang et al. (2022); OpenAI (2023a) with human-level generative capabilities presents tremendous opportunities across diverse domains (Deng et al., 2023; Li et al., 2024; Wang et al., 2023). However, their ability to synthesize high-quality text also raises widespread concerns about potential misuse, including the dissemination of misinformation  Zellers et al. (2019); Chen et al. (2023a) and the facilitation of academic dishonesty  Stokel-Walker (2022). This necessitates developing techniques to reliably attribute generated text to AI systems.
@@
Existing approaches typically fall into two main paradigms. The first type attempts to distinguish machine-generated text by hunting for inductive statistical or linguistic patterns Gehrmann et al. (2019); Mitchell et al. (2023); Zellers et al. (2019); OpenAI (2023b), employing methods that span from basic manual feature engineering to the intricate training of complex classifiers. However, as generative models continue improving, their outputs increasingly resemble the pattern of human writing, rendering statistical detectors ineffective Dou et al. (2022); Sadasivan et al. (2023); Chakraborty et al. (2023). The second paradigm promotes a more proactive approach, advocating for direct intervention in the generative process to actively watermark model outputs Kirchenbauer et al. (2023); Christ et al. (2023); Zhao et al. (2023). This strategy embeds identifiable fingerprints within machine-generated text, enabling provenance verification. As LLMs’ capabilities continue to grow, this approach is more effective in detecting LLM-generated text Sadasivan et al. (2023). However, introducing watermarks during text generation can significantly impact output quality, which has been a consistent challenge for model developers - how to effectively watermark while preserving text quality.
@@
In this paper, we introduce a novel approach to text watermarking by leveraging engineered lexical redundancy during the decoding phase of language generation. Our method utilizes the comprehensive set of tokens available to a language model, clustering them based on overlapping semantic or syntactic functionalities to create sets of interchangeable tokens. This process simulates redundancy within the lexical space, akin to the surplus pixels in images that facilitate watermarking in multimodal data Nikolaidis and Pitas (1999); Samuel and Penzhorn (2004). The motivation for this strategy arises from the challenge of applying traditional watermarking techniques to textual data. In contrast to the inherent redundancy found in images, the discrete and succinct nature of textual data offers little to no native redundancy, making it challenging to exploit redundancy in the textual space Zhou et al. (2021); He et al. (2022). By engineering lexical redundancy, our method not only surmounts the limitations imposed by the inherent properties of natural language but also paves the way for secure and efficient text watermarking.
@@ 2 Related Work
Early works on AI-generated text detection develop post-hoc detection methods to analyze machine-generated text by treating the problem as a binary classification task OpenAI (2019); Jawahar et al. (2020); Mitchell et al. (2023). For instance, OpenAI has fine-tuned RoBERTa Liu et al. (2019) to distinguish between human and GPT-2 generated texts OpenAI (2019). However, existing detectors are found to be fragile against adversarial attacks Wolff (2020) and biased towards non-native English writers Liang et al. (2023). Moreover, as LLMs continue to advance, their generated outputs more closely resemble human-written text, rendering these methods progressively less effective.
On the other side, watermarking, traditionally a copyright marking method Adi et al. (2018); Rouhani et al. (2018), involves developers, users, and regulatory entities. Developers choose an algorithm to subtly embed hidden modifications into data, which can be altered during user transmission. Regulatory bodies can later extract this information to trace and regulate AI-generated content Atallah et al. (2001); Wilson et al. (2014); Hacker et al. (2023). In the context of natural languages, watermarking typically involves modifying content or structure. For example, rule-based methods Stefan et al. (2000) or carefully designed neural encoders Yang et al. (2022); Ueoka et al. (2021) encrypt messages into text, which are then extracted using the corresponding rules and neural decoder. The discrete nature of natural language, however, presents a considerable challenge to this approach, as modifications can unintentionally degrade text quality or alter its intended meaning.
For the detection of LLM-generated texts, a pioneering watermarking technique Kirchenbauer et al. (2023) partitions tokens into ’green’ and ’red’ lists, biases output distribution towards ’green’ tokens, and creates patterns that are detectable yet imperceptible to humans. Nevertheless, while yielding promising detection results, these methods may still degrade the textual quality and be vulnerable to the paraphrase attack. Current efforts Christ et al. (2023); Fernandez et al. (2023); Zhao et al. (2023) in this field aim to develop more robust watermarking methods capable of defending various user attacks.
@@ 3 Method
In this section, we begin by providing a summary of the preliminaries related to text watermarking. Subsequently, we delve into an investigation of redundancy in the lexical space and demonstrate how this redundancy can be leveraged to develop a watermarking algorithm that achieves a higher degree of losslessness for large language models. Finally, we employ mathematical analysis to elucidate the benefits of our proposed method.
@@ 3.3 WatME: Exploit the Lexical Redundancy
We then detail the WatME encoding process, outlined in Alg. 1, which employs a two-step partitioning process to form a green and red list. The first partition occurs within the redundant lexical clusters $C$ that we have identified within $S$, while the second takes place among the remaining part in the vocabulary denoted as $\mathcal{V}\setminus\mathcal{S}$. We use $\gamma$ to determine the number of tokens from the mined clusters that are allocated to the green list $G_{t}^{\prime}$ in the first partition. The remaining tokens, based on the principle of mutual exclusivity, are assigned to the red team $R_{t}^{\prime}$. The second partition continues to allocate words to the green list $G_{t}$ from the remaining vocabulary until the combined size of the green teams from both steps reaches the predefined limit, $\gamma$. The rest of the process follows the steps outlined in the vanilla watermarking of Alg. 2.
@@ 5.1 Experimental Setups
+ KGW-Mark 5.61-50.0% 0.8886 57.16-39.9% 84.33-9.1% 48.20-45.4% 0.8416 7.00 0.9724
+ Gumbel-Mark 7.28-35.1% 0.9121 45.90-51.7% 92.78-0.0% 42.59-51.7% 0.4931 39.93 0.9422
+ Unbiased-Mark 10.24-8.7% 0.5478 44.06-53.7% 93.76+1.1% 41.43-53.0% 0.5051 15.62 0.5451
+ Provable-Mark 5.16-54.01% 0.9052 64.14-32.6% 91.68-1.2% 58.80-33.4% 0.9555 10.21 0.9623
+ WatMEdictionary 9.17-18.3% 0.8995 69.28-27.2% 88.25-4.9% 61.14-30.7% 0.8848 5.32 0.9804
+ WatMEprompting 5.84-48.0% 0.9128 55.83-41.3% 95.10+2.5% 50.39-42.9% 0.8659 6.89 0.9724
@@
+ KGW-Mark 13.87-20.8% 0.7870 74.05-21.1% 87.52+0.3% 64.81-20.1% 0.7417 11.62 0.9679
+ Gumbel-Mark 9.02-48.5% 0.7077 68.30-27.2% 87.27-0.0% 59.61-27.2% 0.4647 48.93 0.8617
+ Unbiased-Mark 17.89+2.2% 0.5508 70.38-25.0% 88.86+1.8% 62.54-23.7% 0.4855 19.93 0.5000
+ Provable-Mark 12.21-30.27% 0.8020 74.42-20.7% 96.70+10.8% 71.96-12.2% 0.8796 10.21 0.9582
+ WatMEdictionary 14.78-15.6% 0.8044 78.95-15.9% 97.43+11.6% 76.92-6.1% 0.7897 10.96 0.9582
+ WatMEprompting 16.22-7.4% 0.7843 69.65-25.8% 97.45-11.5% 67.87-17.2% 0.7396 11.54 0.9519
@@
We compared our model with four established baselines. First, KGW-Mark (Vanilla watermarking) Kirchenbauer et al. (2023), which categorizes teams into ’red’ and ’green’ to facilitate detection. Second, Gumbel-Mark Kuditipudi et al. (2023), which employs a Gumbel-Softmax distribution to introduce stochasticity into the watermarking process. Third, Unbiased-Mark Hu et al. (2024), which implements reweighting techniques to maintain the expected output distribution of the LLM during watermarking. Lastly, Provable-Mark Zhao et al. (2023), which uses a fixed hash key during watermarking to achieve provably better performance.
@@ 5.2 Main Results
Despite the trade-off between text quality and detection performance, WatME’s detection efficacy matched that of the Vanilla Watermark while also enhancing model capabilities, as evidenced by similar AUROC scores—suggesting our algorithm attained a better equilibrium than the baseline. In contrast, the Gumbel-Mark method noticeably compromised detection performance, particularly in aligned models and when generating short responses (TruthfulQA). Additionally, more performance results under different watermark strengths are presented in Discussion 6.3.
@@ 6.1 Analysis of Clustering Methods
To analyse redundant clusters from diverse methods, we set evaluation criteria to ensure analytical rigour. These criteria spanned semantic consistency, contextual appropriateness, and grammatical consistency, which are essential aspects of cluster quality. Two annotators used a rating scale of 0, 1, 2 to annotate the clusters. A score of ’2’ indicated high or ideal consistency, ’1’ denoted moderate or usable consistency, and ’0’ identified low or unusable consistency within a cluster. The kappa value for the annotations is 0.657. Figure 2(a) shows both methods met usability, but fell short of ideal effectiveness. The dictionary approach was superior in semantic coherence due to its utilization of lexical databases. Conversely, the prompting method outperformed in contextual and grammatical consistency, reflecting the varied linguistic corpus training of LLMs. This suggests the potential benefits of a combined approach, a topic reserved for future research.
@@ 6.2 Robustness Against Attacks
In addition to affecting the performance of LLMs, watermarks are also vulnerable to attacks aimed at their removal. To evaluate the robustness of our method, we conducted tests against two prevalent types of attacks: substitution attacks and paraphrase attacks. For the substitution attack, we evaluated 200 examples from GSM8k, with various token replacement ratios. As shown in Figure 2(b), WatME consistently outperformed the baseline method in the robustness of detection across different levels of token replacement. For paraphrase attacks, we use a powerful paraphraser, llama-2-chat-13B, to extensively rewrite the watermarked text generated by llama-2-7b. We provided it with the prompt: "Please paraphrase the following text, altering the wording significantly yet preserving the original meaning and length." We then subjected our system to these rewritten samples using 200 entries from both GSM8k and TruthfulQA. The results are presented in Tables 2.
We offer two perspectives to understand the robustness of WatME: (1) Intuitively, for substitution attacks, the effect on watermarking depends on whether it triggers a token swap between the ’red’ and ’green’ teams: a swap affects the detection, while no swap means the watermark remains intact. With KGW-Mark, semantically similar tokens may be allocated to one team, resulting in a substitution invariably causing a swap. In contrast, WatME is intentionally designed to prevent this scenario. Therefore, the likelihood of a red-green swap—and consequently the impact on the watermark—is reduced in WatME compared to KGW. (2) From an encryption viewpoint, whereas KGW-Mark relies on a single division between teams, WatME employs multiple divisions—the number of clusters plus one (|C|+1), as outlined in Algorithm 2. Though these multiple partitions are computationally equivalent to a single partition due to efficient parallel matrix operations (explained in Appendix D), they introduce a higher level of complexity and robustness to the encryption process.
Method Dataset Original Para. Attack
KGW-Mark GSM8k 0.885 0.745
WatME 0.955 0.910
KGW-Mark TruthfulQA 0.924 0.528
WatME 0.949 0.673
Table 2: Detection robustness against paraphrasing attacks.
@@ 6.3 Performance Trade-offs at different Delta
The efficacy of the Watermark is influenced by the hyperparameter, Delta, which controls the watermark strength. An increase in Delta facilitates easier watermark detection but at the cost of severer impact on the LLMs. We analyse the TruthfulQA and GSM8K datasets. Figure 3 shows WatME consistently achieved a more favourable balance between watermark robustness and LLM performance across various Delta settings, surpassing Vanilla Watermark. Notably, the performance curves of WatME are strictly better than that of Vanilla, indicating that at equivalent watermark strengths, WatME always maintains superior performance compared to Vanilla Watermark.
@@ 6.4 Aligned vs Unaligned Models
Our examination of the response sensitivity to watermarking in aligned and unaligned models involved analyzing their output distributions on the TruthfulQA and GSM8K datasets. We computed the average entropy for tokens in the generated answers and found that aligned models exhibit markedly lower entropy, suggesting more deterministic response patterns, as illustrated in Figure 4. This pronounced certainty in aligned models’ outputs presents a challenge for watermarking because of the limited variability that is essential for effective watermark encoding.
@@ 7 Conclusion
Lastly, our LLMs-based cluster generation approach is influenced by the robustness of the prompting methods. Different prompt constructions can lead to varying outcomes Zhao et al. (2021); Chen et al. (2023b, 2024), represents a limitation that warrants further discussion and exploration in future work.
@@ Appendix F Examples of Redundant Clusters
We present some examples of mined clusters at 3.
Dictionary-based Method LLM-based Method
’should’, ’must’, ’would’ ’must’, ’ought’, ’should’
’job’, ’pursuit’, ’operation’, ’profession’, ’career’, ’employment’, ’behavior’ ’job’, ’task’, ’work’
’inside’, ’in’ ’_inside’, ’_inner’, ’_within’
Table 3: Examples of Redundant Clusters.
Method ROUGE-L AUROC
ChatGLM 3-6b 11.29 -
+KGW-Mark 8.89 0.8415
+WatME_prompting 10.23 0.8514
Table 4: CLTS Results
Appendix G Multilingual Performance Testing
We expand our evaluation to include the Chinese Long Text Summarization Dataset (CLTS) and a bilingual Large Language Model (LLM), ChatGLM3-6b. This model employs Byte Pair Encoding (BPE) tokenization with a vocabulary size of 65k—double that of the Llama 2 model which has a 32k vocabulary size. Synonym mining, a critical step in our process, was conducted using the ChatGLM3-13B model. The performance of different watermarking methods was evaluated using the ROUGE-L and AUROC metrics, as shown in Table 4. The results highlight that watermarking with WatME considerably enhances detection robustness compared to the baseline method, maintaining effectiveness despite varying levels of token replacement. This improvement underscores WatME’s capability to integrate seamlessly without compromising the natural language generation quality.